# Confidence on the Focal: Conformal Prediction with Selection-Conditional Coverage

### v1

@@ B.7 Proof of Proposition 6
    by the choice of $T^{\textnormal{swap}(i,j)}(y)$. The above implies that $T^{\textnormal{swap}(i,j)}(y)\geq T_{j}^{\textnormal{fixed}}(0,0)\Rightarrow T^{\textnormal{swap}(i,j)}=T_{j}^{\textnormal{fixed}}(0,0)$. Consequently, we have
    $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)=\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{fixed}}(0,0)\}\in\mathfrak{S}.$
    We have thus shown that $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$.
  • $Y_{i}>c_{i}$.
    Since $y>c_{n+j}$ and $Y_{i}>c_{i}$, for any $t\in{\mathcal{T}}$ there is
    $1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq t,Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq t,y\leq c_{n+j}\}=1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{\ell}\geq t,Y_{\ell}\leq c_{\ell}\}.$
    implying that $T_{j}^{\textnormal{fixed}}(0,1)=T^{\textnormal{swap}(i,j)}(y)$. Consequently, $\{\widehat{S}_{i}\geq T_{j}^{\textnormal{fixed}}(0,1)\}=\{\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y)\}$ and
    $\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{fixed}}(0,1)\}=\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y).$
    Then there is $i\in\widehat{{\mathcal{R}}}^{\textnormal{fixed}}_{n+j}(y)\iff i\in\widehat{{\mathcal{R}}}_{n+j}(y)$.
Combining the cases of $Y_{i}\leq c_{i}$ and $Y_{i}>c_{i}$, we have ^$\widehat{{\mathcal{R}}}^{\textnormal{fixed}}_{n+j}=\widehat{{\mathcal{R}}}_{n+j}(y)$, and therefore ^$\widehat{{\mathcal{R}}}^{\textnormal{fixed}}_{n+j}(y)=\widehat{{\mathcal{R}}}_{n+j}(y)$, for any $y>c_{n+j}$. The proof is now complete.
B.8 Proof of Proposition 7
We start from making the observation that $\widehat{{\mathcal{S}}}_{\textnormal{BH}}=\{j\in[m]:\widehat{S}_{n+j}\geq T^{\textnormal{BH}}\}$, where
$T^{\textnormal{BH}}=\inf\bigg{\{}t\in{\mathcal{T}}:\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\{\widehat{S}_{i}\geq t,Y_{i}\leq c_{i}\}}{(\sum_{j\in{\mathcal{I}}_{\textnormal{test}}}\mathds{1}\{\widehat{S}_{j}\geq t\})\vee 1}\frac{m}{1+n}\leq q\bigg{\}}.$
To see this, let $j^{*}=\sum^{m}_{j=1}\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{BH}}\}$ and $\widehat{S}_{n+(1)},\widehat{S}_{n+(2)},\ldots,\widehat{S}_{n+(m)}$ denote the order statistics of $\{\widehat{S}_{n+j}\}_{j\in[m]}$ in an decending order. It then suffices to show that $j^{*}=k^{*}$. When $k^{*}\geq 1$,
$\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\{\widehat{S}_{i}\geq\widehat{S}_{n+(k^{*})},Y_{i}\leq c_{i}\}}{(\sum_{j\in{\mathcal{I}}_{\textnormal{test}}}\mathds{1}\{\widehat{S}_{j}\geq\widehat{S}_{n+(k^{*})}\})\vee 1}\frac{m}{1+n}=p_{(k^{*})}\frac{m}{k^{*}}\leq q,$ (48)
$p_{(j^{*})}\frac{m}{j^{*}}$ =$\overset{\rm(a)}{=}\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq\widehat{S}_{n+(j^{*})},Y_{i}\leq c_{i}\}}{n+1}\cdot\frac{m}{\sum^{m}_{j=1}\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{BH}}\}}$
$\overset{\rm(b)}{\leq}\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq T^{\textnormal{BH}},Y_{i}\leq c_{i}\}}{\sum^{m}_{j=1}\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{BH}}\}}\cdot\frac{m}{n+1}\overset{\rm(c)}{\leq}q.$
$\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq t,Y_{i}\leq c_{i}\}}{(\sum^{m}_{j=1}\mathds{1}\{\widehat{S}_{n+j}\geq t\})\vee 1}\frac{m}{n+1}=\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq t,Y_{i}\leq c_{i}\}}{j(t)}\frac{m}{n+1}$
[…]
where the last step is because $k^{*}=0$. The above implies that $T^{\textnormal{BH}}>\widehat{S}_{n+(1)}\Rightarrow j^{*}=0$. The equivalence is therefore established.
As before, we define for any $i\in[n]$, $j\in[m]$, and $y\in{\mathcal{Y}}$ the swapped stopping time:
$T^{{\textnormal{swap}(i,j)}}(y)=\inf\bigg{\{}t\in{\mathcal{T}}:1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq t,Y_{i}\leq c_{i}\}+\mathds{1}\{\widehat{S}_{n+j}\geq t,y\leq c_{n+j}\}\leq q(n+1)\bigg{\}}.$
Then $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)=\{\ell\in[m]:\widehat{S}^{\textnormal{swap}(i,j)}_{n+\ell}\geq T^{\textnormal{swap}(i,j)}(y)\}$.
We are now ready to prove the proposition. We consider the two cases of $y\leq c_{n+j}$ and $y>c_{n+j}$ separately.
Case 1: $y\leq c_{n+j}$.
We discuss the cases of $Y_{i}\leq c_{i}$ and $Y_{i}>c_{i}$.
  • $Y_{i}\leq c_{i}$. Suppose $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$. Then $\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y)$ and $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)\in\mathfrak{S}$. Since $\mathds{1}\{\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y),T_{i}\leq c_{i}\}=1$ and $y\leq c_{n+j}$,
    $\frac{\sum_{\ell\in[n]}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{swap}(i,j)}(y),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{swap}(i,j)}(y)\}}{\sum_{\ell\in[m]}\mathds{1}\{\widehat{S}_{n+\ell}\geq T^{\textnormal{swap}(i,j)}(y)\}}\frac{m}{1+n}$
[…]
    where the last step follows from the definition of $T^{\textnormal{swap}(i,j)}(y)$. By the choice of $T^{\textnormal{BH}}_{j}(1,0)$, there is $T^{\textnormal{swap}(i,j)}(y)\geq T^{\textnormal{BH}}_{j}(1,0)$. As a result, $\widehat{S}_{i}\geq T^{\textnormal{BH}}_{j}(1,0)$. Conversely,
    $\frac{1+\sum_{\ell\in[n]\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{BH}}_{j}(1,0),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{BH}}_{j}(1,0),y\leq c_{n+j}\}}{\sum_{\ell\in[m]}\mathds{1}\{\widehat{S}_{n+\ell}\geq T^{\textnormal{BH}}_{j}(1,0)\}}\frac{m}{1+n}$
[…]
    following from the definition of $T^{\textnormal{BH}}_{j}(1,0)$. We therefore have $T^{\textnormal{BH}}_{j}(1,1)\geq T^{\textnormal{swap}(i,j)}(y)$, and then $T^{\textnormal{BH}}(1,0)=T^{\textnormal{swap}(i,j)}(y)$. The above also implies that
    $\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T^{\textnormal{BH}}_{j}(1,0)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T^{\textnormal{swap}(i,j)}(y)\}=\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)\in\mathfrak{S}.$
    By definition, $i\in\widehat{{\mathcal{R}}}^{\textnormal{BH}}_{n+j}(y)$.
    On the other hand, suppose $i\in\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{BH}}(y)$, we have $\widehat{S}_{i}\geq T^{\textnormal{BH}}_{j}(1,0)$ and $\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T^{\textnormal{BH}}_{j}(1,0)\}\in\mathfrak{S}$. We can check that
    $\frac{1+\sum_{\ell\in[n]\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{BH}}_{j}(1,0),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{BH}}_{j}(1,0),y\leq c_{n+j}\}}{\sum_{\ell\in[m]}\mathds{1}\{\widehat{S}_{n+\ell}\geq T^{\textnormal{BH}}_{j}(1,0)\}}\frac{m}{1+n}$
[…]
    by the definition of $T^{\textnormal{BH}}_{j}(1,0)$; the definition of $T^{\textnormal{swap}(i,j)}(y)$ implies that $T^{\textnormal{BH}}_{j}(1,0)\geq T^{\textnormal{swap}(i,j)}(y)$. Conversely, since $\mathds{1}\{\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y),Y_{i}\leq c_{i}\}=1$ and $y\leq c_{n+j}$, there is
    $\frac{\sum_{\ell\in[n]}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{swap}(i,j)}(y),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{swap}(i,j)}(y),y\leq c_{n+j}\}}{\sum_{\ell\in[m]}\mathds{1}\{\widehat{S}_{n+\ell}\geq T^{\textnormal{swap}(i,j)}(y)\}}\frac{m}{n+1}$
[…]
    and therefore $T^{\textnormal{swap}(i,j)}\geq T^{\textnormal{BH}}_{j}(1,0)\Rightarrow T^{\textnormal{swap}(i,j)}(y)=T^{\textnormal{BH}}_{j}(1,0)$. We then have that $\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y)$ and
    ^$\widehat{{\mathcal{R}}}^{\textnormal{swap}(i,j)}(y)=\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T^{\textnormal{swap}(i,j)}(y)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T^{\textnormal{BH}}_{j}(1,0)\}\in\mathfrak{S}.$
    That is, $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$. We have thus shown that $i\in\widehat{{\mathcal{R}}}^{\textnormal{BH}}_{n+j}\iff i\in\widehat{{\mathcal{R}}}_{n+j}(y)$ for any $i\in{\mathcal{I}}_{\textnormal{calib}}$ such that $Y_{i}\leq c_{i}$.
  • If $Y_{i}>c_{i}$, then for any $t\in{\mathcal{T}}$,
    $\frac{1+\sum_{\ell\in[n]\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq t,Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq t,y\leq c_{n+j}\}}{\sum_{\ell\in[m]}\mathds{1}\{\widehat{S}_{n+\ell}\geq t,Y_{\ell}\leq c_{\ell}\}}\frac{m}{1+n}$
[…]
    By their definition, we have $T^{\textnormal{BH}}_{j}(1,1)=T^{\textnormal{swap}(i,j)}$. As a consequence,
    $\widehat{S}_{i}\geq T^{\textnormal{BH}}_{j}(1,1)\iff\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y)\iff j\in\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y).$
    and
    $\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T^{\textnormal{BH}}_{j}(1,1)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T^{\textnormal{swap}(i,j)}(y)\}=\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y).$
    As a result, $i\in\widehat{{\mathcal{R}}}^{\textnormal{BH}}_{n+j}(y)\iff i\in\widehat{{\mathcal{R}}}_{n+j}(y)$ for any $i\in{\mathcal{I}}_{\textnormal{calib}}$ such that $Y_{i}>c_{i}$.
Concluding the above two cases, we have shown that ^$\widehat{{\mathcal{R}}}_{n+j}(y)=\widehat{{\mathcal{R}}}^{\textnormal{BH}}(y)$ when $y\leq c_{n+j}$.
Case 2: $y>c_{n+j}$.
We again prove ^$\widehat{{\mathcal{R}}}_{n+j}(y)=\widehat{{\mathcal{R}}}^{\textnormal{BH}}_{n+j}(y)$ for $y>c_{n+j}$ in two cases.
  • $Y_{i}\leq c_{i}$. Suppose $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$. Then $\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y)$ and $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)\in\mathfrak{S}$. Since $\mathds{1}\{\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y),Y_{i}\leq c_{i}\}=1$ and $y>c_{n+j}$,
    $\frac{\sum_{\ell\in[n]}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{swap}(i,j)}(y),Y_{\ell}\leq c_{\ell}\}}{\sum_{\ell\in[m]}\mathds{1}\{\widehat{S}_{n+\ell}\geq T^{\textnormal{swap}(i,j)}(y)\}}\frac{m}{n+1}$
[…]
    by the construction of $T^{\textnormal{swap}(i,j)}(y)$; therefore, $T^{\textnormal{swap}(i,j)}(y)\geq T^{\textnormal{BH}}_{j}(0,0)$. Meanwhile, since $\mathds{1}\{\widehat{S}_{i}\geq T^{\textnormal{BH}}_{j}(0,0),Y_{i}\leq c_{i}\}=1$,
    $\frac{1+\sum_{\ell\in[n]\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{BH}}_{j}(0,0),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{BH}}_{j}(0,0),y\leq c_{n+j}\}}{\sum_{\ell\in[m]}\mathds{1}\{\widehat{S}_{n+\ell}\geq T^{\textnormal{BH}}_{j}(0,0)\}}\frac{m}{n+1}$
[…]
    following from the definition of $T^{\textnormal{BH}}_{j}(0,0)$. Therefore $T^{\textnormal{BH}}_{j}(0,0)\geq T^{\textnormal{swap}(i,j)}(y)\Rightarrow T^{\textnormal{BH}}_{j}(0,0)=T^{\textnormal{swap}(i,j)}(y)$. As a result, $\widehat{S}_{i}\geq T^{\textnormal{BH}}(0,0)$ and
    $\{\ell\in[m]:\widehat{S}_{n+\ell}\geq T^{\textnormal{BH}}_{j}(0,0)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}\geq T^{\textnormal{swap}(i,j)}(y)\}=\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)\in\mathfrak{S}.$
    It follows from the definition that $i\in\widehat{{\mathcal{R}}}^{\textnormal{BH}}_{n+j}(y)$.
    Conversely, if $i\in\widehat{{\mathcal{R}}}^{\textnormal{BH}}_{n+j}(y)$, then $\widehat{S}_{i}\geq T^{\textnormal{BH}}_{j}(0,0)$ and therefore
    $\frac{1+\sum_{\ell\in[n]\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{BH}}_{j}(0,0),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T^{\textnormal{BH}}_{j}(0,0),y\leq c_{n+j}\}}{\sum_{\ell\in[m]}\mathds{1}\{\widehat{S}_{n+\ell}\geq T^{\textnormal{BH}}_{j}(0,0),Y_{\ell}\leq c_{\ell}\}}\frac{m}{1+n}$
[…]
    which is a result of the definition of $T^{\textnormal{BH}}(0,0)$. The above also implies that $T^{\textnormal{BH}}(0,0)\geq T^{\textnormal{swap}(i,j)}(y)$. Next,
    $\frac{\sum_{\ell\in[n]}\mathds{1}\{\widehat{S}_{\ell}\geq T^{\textnormal{swap}(i,j)}(y),Y_{\ell}\leq c_{\ell}\}}{\sum_{\ell\in[m]}\mathds{1}\{\widehat{S}_{n+\ell}\geq T^{\textnormal{swap}(i,j)}\}}\frac{m}{n+1}$
[…]
    where the last step is a result of $T^{\textnormal{swap}(i,j)}(y)$’s definition. The above implies that $T^{\textnormal{swap}(i,j)}(y)\geq T^{\textnormal{BH}}_{j}(0,0)\Rightarrow T^{\textnormal{BH}}_{j}(0,0)=T^{\textnormal{swap}(i,j)}(y)$. We then have
    $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)=\{\ell\in[m]:\widehat{S}_{n+\ell}\geq T^{\textnormal{swap}(i,j)}(y)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}\geq T^{\textnormal{BH}}_{j}(0,0)\}\in\mathfrak{S}.$
    It follows from the definition that $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$. Combining the above, we have shown that $i\in\widehat{{\mathcal{R}}}^{\textnormal{BH}}_{n+j}(y)\iff i\in\widehat{{\mathcal{R}}}^{\textnormal{swap}(i,j)}(y)$ for any $i$ such that $Y_{i}\leq c_{i}$.
  • $Y_{i}>c_{i}$. In this case, we have for any $t\in{\mathcal{T}}$ that
    $\frac{1+\sum_{\ell\in[n]\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq t,Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq t,y\leq c_{n+j}\}}{\sum_{\ell\in[m]}\mathds{1}\{\widehat{S}_{n+\ell}\geq t,Y_{\ell}\leq c_{\ell}\}}\frac{m}{1+n}$
[…]
    By their definitions, there is $T^{\textnormal{swap}(i,j)}(y)=T^{\textnormal{BH}}_{j}(0,1)$, and therefore
    $\widehat{S}_{i}\geq T^{\textnormal{BH}}(0,1)\iff\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y)$
    and
    $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)=\{\ell\in[m]:\widehat{S}_{n+\ell}\geq T^{\textnormal{swap}(i,j)}(y)\}=\{\ell\in[m]:\widehat{S}_{n+\ell}\geq T^{\textnormal{BH}}_{j}(0,1)\}.$
    It follows from the definition that $i\in\widehat{{\mathcal{R}}}^{\textnormal{BH}}(y)\iff i\in\widehat{{\mathcal{R}}}_{n+j}(y)$ for any $i$ such that $Y_{i}>c_{i}$.
Combining the two cases, we have shown that ^$\widehat{{\mathcal{R}}}^{\textnormal{BH}}(y)=\widehat{{\mathcal{R}}}_{n+j}(y)$ for any $y>c_{n+j}$. The proof is now complete.
B.9 Proof of Proposition 8
Recall that
where ^$\widehat{{\mathcal{R}}}_{n+j}(y)$ ia the reference set. We consider the following three cases: $S(X_{n+j},y)<\eta^{-}$, $S(X_{n+j},y)>\eta^{+}$ and $\eta^{-}\leq S(X_{n+j},y)\leq\eta^{+}$. The case where $\eta^{-}\leq S(X_{n+j},y)\leq\eta^{+}$ is obvious, so we discuess the remaining two cases in the following.
As before, we define $\eta^{\textnormal{swap}(i,j)}(y)$ to be the $K$-th smallest elements in $\{\widehat{S}^{\textnormal{swap}(i,j)}_{\ell}\}_{\ell\in{\mathcal{I}}_{\textnormal{calib}}}$.
Case 1: $S(X_{n+j},y)<\eta^{-}$.
^$\widehat{{\mathcal{R}}}_{n+j}(y)={\widetilde{R}}_{n+j}^{(1)}\,:=\,\{i\in{\mathcal{I}}_{\textnormal{calib}}:S_{i}\leq\eta^{-},{\mathcal{L}}(X_{i},\eta)=1\}\cup\{i\in{\mathcal{I}}_{\textnormal{calib}}:S_{i}>\eta^{-},{\mathcal{L}}(X_{i},\eta^{-})=1\}.$
To see this, note that when $S_{i}\leq\eta^{-}$, we have $\eta^{{\textnormal{swap}(i,j)}}(y)=\eta$. and therefore ${\mathcal{L}}(X_{i},\eta)={\mathcal{L}}(X_{i},\eta^{{\textnormal{swap}(i,j)}}(y))$; when $S_{i}>\eta^{-}$, then $\eta^{{\textnormal{swap}(i,j)}}(y)=\eta^{-}$, and therefore ${\mathcal{L}}(X_{i},\eta^{-})={\mathcal{L}}(X_{i},\eta^{{\textnormal{swap}(i,j)}}(y))$. As a result, we have that ^$\widehat{{\mathcal{R}}}_{n+j}(y)={\widetilde{R}}_{n+j}^{(1)}$ and therefore
$V(X_{n+j},y)\leq{\textnormal{Quantile}}\big{(}\{V_{i}\}_{i\in{\widetilde{R}}_{n+j}^{(1)}}\cup\{\infty\},1-\alpha\big{)}$
Case 2: $S(X_{n+j},y)>\eta^{+}$.
^$\widehat{{\mathcal{R}}}_{n+j}(y)={\widetilde{R}}_{n+j}^{(2)}\,:=\,\{i\in{\mathcal{I}}_{\textnormal{calib}}:S_{i}\leq\eta,{\mathcal{L}}(X_{i},\eta^{+})=1\}\cup\{i\in{\mathcal{I}}_{\textnormal{calib}}:S_{i}>\eta,{\mathcal{L}}(X_{i},\eta)=1\}.$
This is because when $S_{i}\leq\eta$, $\eta^{{\textnormal{swap}(i,j)}}(y)=\eta^{+}$ and therefore ${\mathcal{L}}(X_{i},\eta^{{\textnormal{swap}(i,j)}}(y))={\mathcal{L}}(X_{i},\eta^{+})$; when $S_{i}>\eta$, then $\eta^{{\textnormal{swap}(i,j)}}(y)=\eta$, and therefore ${\mathcal{L}}(X_{i},\eta^{{\textnormal{swap}(i,j)}}(y))={\mathcal{L}}(X_{i},\eta)$. Consequently,
$V(X_{n+j},y)\leq{\textnormal{Quantile}}\big{(}\{V_{i}\}_{i\in{\widetilde{R}}_{n+j}^{(2)}}\cup\{\infty\},1-\alpha\big{)}.$
Combining the three cases, we finish the proof.
Appendix C Auxiliary results
C.1 The conformal p-values
The conformal p-value and its random counterpart defined in Jin and Candès, 2023b are
$\bar{p}_{j}=\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{S_{i}\geq\widehat{S}_{n+j}\}}{n+1}\text{ and }$
$\bar{p}^{\textnormal{rand}}_{j}=\frac{U_{n+j}(1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{S_{i}=\widehat{S}_{n+j}\})+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{S_{i}>\widehat{S}_{n+j}\}}{n+1}.$

### v3

@@ E.1 Connection between selection-conditional coverage and FCR control
$=\frac{\sum_{j=1}^{m}\mathbb{E}[\mathds{1}\{V(X_{n+j},Y_{n+j})>\widehat{q}_{m,n},j\in\widehat{\mathcal{S}}\}]}{\sum_{j=1}^{m}\mathbb{E}[\mathds{1}\{j\in\widehat{\mathcal{S}}\}]}$
For the numerator, we have
$\Big{|}\frac{1}{m}\sum_{j=1}^{m}\mathbb{E}\big{[}\mathds{1}\{V_{n+j}>\widehat{q}_{m,n},j\in\widehat{\mathcal{S}}\}\big{]}-\frac{1}{m}\sum_{j=1}^{m}\mathbb{E}\big{[}\mathds{1}\{V_{n+j}>\widehat{q}_{m,n},j\in\widehat{\mathcal{S}}^{*}\}\big{]}\Big{|}$
$\leq\frac{1}{m}\mathbb{E}\big{[}|\widehat{{\mathcal{S}}}\Delta\widehat{{\mathcal{S}}}^{*}|\big{]}=\mathbb{E}\bigg{[}\frac{|\widehat{{\mathcal{S}}}\Delta\widehat{{\mathcal{S}}}^{*}|}{|\widehat{{\mathcal{S}}}^{*}|}\frac{|\widehat{{\mathcal{S}}}^{*}|}{m}\bigg{]}\to 0$
where the last step is because $|\widehat{\mathcal{S}}\Delta\widehat{\mathcal{S}}^{*}|/|\widehat{\mathcal{S}}|{\to}0$ in probability and the dominated convergence theorem. Using (48), we then have
$\frac{1}{m}\sum_{j=1}^{m}\mathbb{E}\big{[}\mathds{1}\{V_{n+j}>\widehat{q}_{m,n},j\in\widehat{\mathcal{S}}\}\big{]}\to G(q^{*}).$
Similarly, for the denominator,
$\frac{1}{m}\Big{|}\sum_{j=1}^{m}\mathbb{E}[\mathds{1}\{j\in\widehat{\mathcal{S}}\}]-\frac{1}{m}\sum_{j=1}^{m}\mathbb{E}[\mathds{1}\{j\in\widehat{{\mathcal{S}}}^{*}\}]\Big{|}\leq\frac{1}{m}\mathbb{E}\bigg{[}\frac{|\widehat{{\mathcal{S}}}\Delta\widehat{{\mathcal{S}}}^{*}|}{|\widehat{{\mathcal{S}}}^{*}|}|\widehat{{\mathcal{S}}}^{*}|\bigg{]}\to 0.$
Therefore, we obtain the same expression for the asymptotic selection-conditional coverage. ∎
E.2 The conformal p-values
Recall that in Section 4.2 of the main text, we introduce the conformal p-value for each $j\in[m]$ as
$p_{j}=\frac{1+\sum_{i\in[n]}\mathds{1}\{Y_{i}\leq c_{i},\widehat{S}_{i}\geq\widehat{S}_{n+j}\}}{n+1},$
where $\widehat{S}_{i}=S(X_{i},c_{i})$, and $S\colon{\mathcal{X}}\times\mathbb{R}\to\mathbb{R}$ is the pre-specified score function. We can construct sharper p-values by introducing extra randomness. Letting $U_{n+1},U_{n+2},\ldots,U_{n+m}$ be i.i.d. uniform random variables on $[0,1]$ that are independent of everything else, we can construct the randomized conformal p-values as
$p_{j}^{\textnormal{rand}}=\frac{U_{n+j}(1+\sum_{i\in[n]}\mathds{1}\{\widehat{S}_{i}=\widehat{S}_{n+j},Y_{i}\leq c_{i}\}+\sum_{i\in[n]}\mathds{1}\{\widehat{S}_{i}>\widehat{S}_{n+j},Y_{i}\leq c_{i}\})}{n+1}.$
The conformal p-values and their random counterparts defined in Jin and Candès, 2023b are
$\bar{p}_{j}=\frac{1+\sum_{i\in[n]}\mathds{1}\{S_{i}\geq\widehat{S}_{n+j}\}}{n+1}\text{ and }$
$\bar{p}^{\textnormal{rand}}_{j}=\frac{U_{n+j}(1+\sum_{i\in[n]}\mathds{1}\{S_{i}=\widehat{S}_{n+j}\})+\sum_{i\in[n]}\mathds{1}\{S_{i}>\widehat{S}_{n+j}\}}{n+1}.$
The validity of $p_{j}$ and $p_{j}^{\textnormal{rand}}$, as well as their relation to $\bar{p}_{j}$ and $\bar{p}_{j}^{\textnormal{rand}}$, is established in the following lemma.